# Neutrino Decoherence via Modified Dispersion

### v1

@@ 1 Introduction
One of the long-standing challenges in fundamental physics is the reconciliation of quantum theory and gravitation. The behaviour of space-time at the tiniest scales remains one of the mysteries as to whether it is susceptible to fluctuations rooted in quantum gravity (QG) and other potential theories. Neutrino, an elementary particle in the Standard Model (SM) of particle physics, provides a great avenue to probe the structure of space-time. The discovery of neutrino oscillation about two decades ago was a watershed moment which established that neutrinos are massive and opened the window to a plethora of new investigations in elementary particle physics. The past two decades have witnessed tremendous experimental progress to develop a clearer picture of neutrino oscillation [1, 2] as well as the measurement of the associated parameters with great accuracy [3, 4, 5]. Current and forthcoming experiments are expected to shed light on a number of unanswered questions, such as exact neutrino mass and its ordering, the octant of the atmospheric mixing angle, and a possible CP violation in the neutrino sector, and to determine all oscillation parameters with extreme precision [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
@@
This work explores the possibility of neutrino and environment interaction in the vacuum due to quantum gravity effects, considering the fluctuating space-time as the stochastic background. Near the Planck scale, the fluctuations in the space-time metric would alter the conventional rules of quantum mechanics for the propagating neutrinos. As a result, uncertainties in neutrino energies and momentum will creep in. These uncertainties would modify the usual dispersion relation of the neutrino system, leading to one possible mechanism of coherence loss. We also compare the decoherence signatures with the current constraints and discuss the prospects of testing it in the forthcoming experiments. It is worth mentioning that neutrino propagation through matter effects also lead to significant changes in the oscillation probabilities. However, we mainly focus on the neutrinos originating from far away astrophysical sources and propagating astronomical distances through vacuum and hence do not consider matter effects.
This article is organized as follows. In Sec. 2 we briefly introduce neutrino oscillation to set our notations and conventions. In Sec. 3, we present the theoretical formalism of OQS and how neutrinos can be treated in this framework. We discuss in Sec. 4 the phenomenological models and derive the modified master equation. Furthermore, we comment on energy dependence of decoherence parameters in Sec. 5. Afterwards, we discuss our results in Sec. 6. In addition, we briefly study $3+1$ flavour scenario in Sec. 7. We finally summarise and conclude in Sec. 8.
@@ 2 A quick look at Neutrino Oscillation
where $\Delta_{ij}=\Delta m^{2}_{ij}/2E$ and $H_{f}=\tilde{U}H_{m}\tilde{U}^{-1}$ is the Hamiltonian matrix in flavour basis and $H_{m}$ is the diagonal Hamiltonian that governs the time evolution of neutrino mass eigenstates with $\tilde{U}$ is unitary matrix. The energies of three different flavour neutrinos are assumed to be same here. In three three-flavour neutrino oscillation paradigm, the probability of oscillation between different neutrino flavours, $\nu_{\alpha}$ and $\nu_{\beta}$, is given by,
@@
Here, $\delta_{\alpha\beta}$ is the Kronecker delta, $L$ is the neutrino path length, and $E$ is the neutrino energy. The oscillation probability for the $\alpha=\beta$ case is known as survival probability, and $\alpha\neq\beta$ is termed as transition probability. Since in practice a neutrino wave is neither detected nor produced with sharp energy or with well defined propagation length, we have to average over the $L/E$ dependence and other uncertainties in the detection and emission of the neutrino wave. Here, we use the Gaussian averaging scheme in which the averaged transition probability takes the form,
@@
where $\sigma$ is the damping parameter responsible for damping the neutrino oscillation probabilities. The damping parameter can be parametrized by the decoherence parameter (described in later sections) following the equivalence established between the two as in Ref [46] [^1] It is to be noted here that the left hand sides of both Eqs. (2.7) and (2.8) denote the same quantity. However, in order to avoid confusion, we do not use delimited-⟨$\langle$..$\rangle$ … ⟩ in the left hand side of Eq. (2.8)., which will damp the probability as,
@@
where $\Gamma_{ab}$ are the decoherence parameters which in turn will depend on the decoherence effects under consideration.
@@ 3 Neutrino as an Open Quantum System
where $\rho$ and $H_{m}$ are the density matrix and the Hamiltonian of the neutrino subsystem, respectively. To incorporate decoherence effects, an additional term $D[\rho]$ is introduced in the Liouville equation (3.2) for the neutrinos, which will allow transitions from pure to mixed quantum mechanical states, is given by,
@@
$D[\rho]=\frac{1}{2}\sum_{k=1}^{N^{2}-1}\left([V_{k},\rho V_{k}^{\dagger}]+[V_{k}\rho,V_{k}^{\dagger}]\right),$ (3.4)
where $V_{k}$, with $k=N^{2}-1$, are a set of operators called as Lindblad operators such that $\text{Tr}(V_{k})=0$ for any $k$, and $\text{Tr}(V_{k}^{\dagger}V_{j})=\delta_{kj}$. For two and three-flavour neutrino mixing cases, $V_{k}$s are represented by the Pauli and Gell-Mann matrices, respectively. Assuming $V_{k}$’s to be Hermitian ($V_{k}^{\dagger}=V_{k}$) and also commuting with the Hamiltonian ($[H_{m},V_{k}]=0$), the dissipator takes the form:
$D[\rho]=\sum_{k=1}^{n}[V_{k},[V_{k},\rho]]=\sum_{k=1}^{n}(\rho V_{k}^{2}+V_{k}^{2}\rho-V_{k}\rho V_{k}),$ (3.5)
@@
$V_{k}$ $=v_{k}{\lambda}_{k},\quad k=1,2,\ldots,8,$ (3.8)
where $\lambda_{k}$ represents the Gell-Mann matrices that satisfies $\lambda_{k}^{\dagger}=\lambda_{k}$, $[\lambda_{a},\lambda_{b}]=2if_{abc}\lambda_{c}$, with structure constants $f_{abc}=-\frac{i}{4}\text{Tr}(\lambda_{a}[\lambda_{b},\lambda_{c}])$ with $\lambda_{0}$ being the $3\times 3$ identity matrix. Substituting these into the GKSL master Eq. (3.3) we obtain:
@@
$D_{\mu\nu}=\begin{pmatrix}-\Gamma_{00}&\beta_{01}&\beta_{02}&\beta_{03}&\beta_{04}&\beta_{05}&\beta_{06}&\beta_{07}&\beta_{08}\\ \beta_{01}&-\Gamma_{11}&\beta_{12}&\beta_{13}&\beta_{14}&\beta_{15}&\beta_{16}&\beta_{17}&\beta_{18}\\ \beta_{02}&\beta_{12}&-\Gamma_{22}&\beta_{23}&\beta_{24}&\beta_{25}&\beta_{26}&\beta_{27}&\beta_{28}\\ \beta_{03}&\beta_{13}&\beta_{23}&-\Gamma_{33}&\beta_{34}&\beta_{35}&\beta_{36}&\beta_{37}&\beta_{38}\\ \beta_{04}&\beta_{14}&\beta_{24}&\beta_{34}&-\Gamma_{44}&\beta_{45}&\beta_{46}&\beta_{47}&\beta_{48}\\ \beta_{05}&\beta_{15}&\beta_{25}&\beta_{35}&\beta_{45}&-\Gamma_{55}&\beta_{56}&\beta_{57}&\beta_{58}\\ \beta_{06}&\beta_{16}&\beta_{26}&\beta_{36}&\beta_{46}&\beta_{56}&-\Gamma_{66}&\beta_{67}&\beta_{68}\\ \beta_{07}&\beta_{17}&\beta_{27}&\beta_{37}&\beta_{47}&\beta_{57}&\beta_{67}&-\Gamma_{77}&\beta_{78}\\ \beta_{08}&\beta_{18}&\beta_{28}&\beta_{38}&\beta_{48}&\beta_{58}&\beta_{68}&\beta_{78}&-\Gamma_{88}\end{pmatrix}$ (3.10)
@@
In theory, the possibility of energy non-conservation within the neutrino subsystem exists, although the energy is inherently conserved for the entire system. However, we strictly adhere to the energy conservation condition within the neutrino subsystem for our analysis. This choice allows for a significant simplification of the dissipator, ultimately resulting in a diagonal form as,
@@ 4.2 MDR-Induced Master Equation
To solve this equation for the density matrix, we move to the interaction picture, where the eigenstate which was previously in the Schrödinger picture, now takes the form,
@@
Note that from here onward the prime represents the respective quantities in the interaction picture. A precise solution to this equation can be obtained by integrating both sides over the interval $[0,t]$, yielding,
[…]
@@
[…]
@@
[…]
@@
[…]
At this stage, we note that Quantum Gravity (QG) theories must include space-time fluctuations at the Planck scale. This can be done by introducing a fluctuating deformation parameter to address the fluctuating minimum length scale, requiring a stochastic treatment to capture the complexities involved. By averaging the fluctuations and introducing the shorthand notation $\langle\rho^{\prime}\rangle=\rho^{\prime}$, the derivative will take the form as,
$\langle\dot{\rho}^{\prime}(t)\rangle=\dot{\rho}^{\prime}(t),$ (4.19)
and Eq. (4.18) takes the form,
[…]
Since $\rho^{\prime}(0)$ is constant, the first term on the r.h.s. of Eq. (4.20) vanishes. Substituting $H_{\mathrm{I}}$ in terms of $H_{\mathrm{0}}$ from Eq. (4.8) in Eq. (4.20) in the second term, we get,
$\dot{\rho}^{\prime}(t)=-\kappa\left[H^{\prime}_{0},\left[H^{\prime}_{0},\rho^{\prime}(t)\right]\right],$ (4.21)
@@
This is the new master equation, which incorporates the decoherence effects arising because of the modified dispersion relation. The first term on r.h.s of (4.23) accounts for trivial unitary dynamics of neutrino oscillations while the second term governs for neutrino-environment interactions due to MDR [^3] In the second term of the r.h.s of Eq. (4.23), the variable $t$ will denote the propagation time of the neutrinos. Since in this work we are mostly dealing with ultra-relativistic neutrinos, the propagation time $t$ can be replaced with the propagation length $L$ as shown in Eq. (6.1).. Furthermore, it can be seen from Eq. (4.23) that it takes different forms depending on the value of $a$, which is an input from the specific theory of quantum gravity that causes the modification to the dispersion relation, which we remain agnostic about. Later we are going to consider a few specific values of $a$ and compare their effects.

### v2

@@ 1 Introduction
One of the long-standing challenges in fundamental physics is the reconciliation of quantum theory and gravitation. The behaviour of space-time at the tiniest scales remains one of the mysteries as to whether it is susceptible to fluctuations rooted in quantum gravity (QG) and other potential theories. Neutrino, an elementary particle in the Standard Model (SM) of particle physics, provides a great avenue to probe the structure of space-time. The discovery of neutrino oscillation about two decades ago was a watershed moment which established that neutrinos are massive and opened the window to a plethora of new investigations in elementary particle physics. The past two decades have witnessed tremendous experimental progress to develop a clearer picture of neutrino oscillation [1, 2] as well as the measurement of the associated parameters with great accuracy [3, 4, 5]. Current and forthcoming neutrino detectors and experiments are expected to shed light on a number of unanswered questions, such as exact neutrino mass and its ordering, the octant of the atmospheric mixing angle, and a possible CP violation in the neutrino sector, and to determine all oscillation parameters with extreme precision [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
@@
This work explores the possibility of neutrino and environment interaction in the vacuum due to quantum gravity effects, considering the fluctuating space-time as the stochastic background. Near the Planck scale, the fluctuations in the space-time metric would alter the conventional rules of quantum mechanics for the propagating neutrinos. As a result, uncertainties in neutrino energies and momentum will creep in. These uncertainties would modify the usual dispersion relation of the neutrino system, leading to one possible mechanism of coherence loss. We also compare the decoherence signatures with the current constraints and discuss the prospects of testing it in the forthcoming experiments. It is worth mentioning that neutrino propagation through matter effects also lead to significant changes in the oscillation probabilities. However, we mainly focus on the neutrinos originating from far away astrophysical sources and propagating astronomical distances through vacuum and hence do not consider matter effects. In this study, we focus on the new physics effects arising from neutrino and environment interactions at high energies, deliberately omitting the decoherence induced by wave-packet separation relevant at lower energies.
This article is organized as follows. In Sec. 2, we briefly introduce neutrino oscillation to set our notations and conventions. In Sec. 3, we present the theoretical formalism of OQS and how neutrinos can be treated in this framework. We discuss in Sec. 4, the phenomenological models and derive the modified master equation. Furthermore, we comment on energy dependence of decoherence parameters in Sec. 5. Afterwards, we discuss our results in Sec. 6. In addition, we briefly study $3+1$ flavour scenario in Sec. 7. We finally summarise and conclude in Sec. 8.
@@ 2 A quick look at Neutrino Oscillation
where $\Delta_{ij}=\Delta m^{2}_{ij}/2E$ and $H_{f}=\tilde{U}H_{m}\tilde{U}^{-1}$ is the Hamiltonian matrix in flavour basis and $H_{m}$ is the diagonal Hamiltonian that governs the time evolution of neutrino mass eigenstates with $\tilde{U}$ is unitary matrix. The energies of three different flavour neutrinos are assumed to be same here. In three-flavour neutrino oscillation paradigm, the probability of oscillation between different neutrino flavours, $\nu_{\alpha}$ and $\nu_{\beta}$, is given by,
@@
Here, $\delta_{\alpha\beta}$ is the Kronecker delta, $L$ is the neutrino path length, and $E$ is the neutrino energy. The oscillation probability for the $\alpha=\beta$ case is known as survival probability, and $\alpha\neq\beta$ is termed as transition probability. Since in practice neutrinos are neither detected nor produced with sharp energy or with well defined propagation length, we have to average over the $L/E$ dependence and other uncertainties in the detection and emission. When neutrinos have length and energy spread $f(L,E)$, the oscillatory terms in Eq. (2.6) fluctuate rapidly and can average out while they reach detectors on Earth. This highlights the need for Gaussian averaging, as the energy spread over long baselines disrupts phase coherence, causing flavor transitions to become probabilistic instead of oscillatory. The Gaussian averging is defined as,
[…]
where $\ell\equiv\langle x\rangle$ and $\sigma\equiv\sqrt{\langle(x-\langle x\rangle)^{2}\rangle}$ are the expectation value and standard deviation, respectively. By taking the Gaussian average of Eq. (2.6) and using $x\equiv\frac{L}{4E}$, we get,
@@
where $\sigma$ is the damping parameter responsible for damping the neutrino oscillation probabilities which can be parametrized by the decoherence parameter (described in later sections) following the equivalence established between the two as in Ref [46] [^1] It is to be noted here that the left hand sides of both Eqs. (2.8) and (2.9) denote the same quantity. However, in order to avoid confusion, we do not use $\langle$$\rangle$ in the left hand side of Eq. (2.9)., which will damp the probability as,
@@
where $\Gamma_{ab}$ are the decoherence parameters which in turn will depend on the decoherence effects under consideration. For our numerical analysis we use the experimentally obtained best-fit values from the NuFit collaboration [64] given in Tab. 1.
Parameter Normal Ordering (NO) Inverted Ordering (IO)
$\sin^{2}\theta_{12}$ $0.308\pm 0.012$ $0.308\pm 0.012$
$\sin^{2}\theta_{23}$ $0.470^{+0.017}_{-0.013}$ $0.550^{+0.012}_{-0.015}$
$\sin^{2}\theta_{13}$ $02215\pm 0.00057 0.0$$02231\pm 0.00056$
$\Delta m^{2}_{21}$ $10^{-5}~\mathrm{eV}^{2}$] $7.49\pm 0.19$ $7.49\pm 0.19$
$\Delta m^{2}_{3\ell}$ $10^{-3}~\mathrm{eV}^{2}$] $2.513\pm 0.021$ $2.484\pm 0.020$
$\delta_{\rm CP}$ [] $212^{+26}_{-41}$ $274^{+22}_{-25}$
Table 1: Neutrino oscillation parameters from NuFit-6.0 (2024), based on global fit with SK atmospheric data.
Moreover, in this work we will consider just the central values of these parameters and the CP-violating phase taken to be zero.
@@ 3 Neutrino as an Open Quantum System
where $\rho$ and $H_{m}$ are the density matrix and the Hamiltonian of the neutrino subsystem, respectively. At this stage, we emphasize that decoherence occurs simultaneously with the oscillation of neutrinos as they propagate. Rather than halting the oscillation, decoherence modifies the oscillatory patterns. Given that neutrino oscillations are determined by the mass basis Hamiltonian, and since decoherence which is described by a dissipator that is expressed in terms of mass basis Hamiltonian, it follows that the phenomenon of decoherence must also be described by the mass basis Hamiltonian. To incorporate decoherence effects, an additional term $D[\rho]$ is introduced in the Liouville equation (3.2) for the neutrinos, which will allow transitions from pure to mixed quantum mechanical states, is given by,
@@
$D[\rho]=\frac{1}{2}\sum_{k=0}^{N^{2}-1}\left([V_{k},\rho V_{k}^{\dagger}]+[V_{k}\rho,V_{k}^{\dagger}]\right),$ (3.4)
where $V_{k}=\sum_{i=0}^{8}v_{ki}\lambda_{i}$ (with $k=0,1,\ldots,N^{2}-1$ where $N$ is the dimension of Hilbert space and in case of three flavour neutrinos $N=3$) represent the Lindblad operators, which are defined to satisfy the orthonormality and tracelessness conditions, $\mathrm{Tr}(V_{k})=0$ and $\mathrm{Tr}(V_{k}^{\dagger}V_{j})=\delta_{kj}$. In general, these operators may exhibit explicit time dependence, $V_{k}\rightarrow V_{k}(t)$, reflecting the temporal variation of the system-environment interaction. For two and three-flavour neutrino mixing cases, $V_{k}$s are represented by the Pauli and Gell-Mann matrices, respectively. Assuming $V_{k}$’s to be Hermitian ($V_{k}^{\dagger}=V_{k}$) and also commuting with the Hamiltonian ($H_{m},V_{k}]=0$), the dissipator takes the form:
$D[\rho]=\sum_{k=1}[V_{k},[V_{k},\rho]]=\sum_{k}(\rho V_{k}^{2}+V_{k}^{2}\rho-V_{k}\rho V_{k}),$ (3.5)
@@
$V_{k}$ $=\sum_{i=0}^{8}v_{ki}\lambda_{i},\quad k=0,1,\ldots,8,$ (3.8)
where $\lambda_{k}$ represents the Gell-Mann matrices that satisfies $\lambda_{k}^{\dagger}=\lambda_{k}$, $\lambda_{a},\lambda_{b}]=2if_{abc}\lambda_{c}$, with structure constants $f_{abc}=-\frac{i}{4}\text{Tr}(\lambda_{a}[\lambda_{b},\lambda_{c}])$ with $\lambda_{0}$ being the $3\times 3$ identity matrix. To have a complete unitary description of neutrino dynamics, it is essential to satisfy the necessary conditions for a density matrix, including trace preservation and complete positivity. The identity matrix ($\lambda_{0}$) plays a critical role in this framework. Additionally, to effectively describe the three-neutrino quantum system, we require a complete nine-element Hermitian basis $\{\lambda_{0},\lambda_{1},\dots,\lambda_{8}\}$ to expand each observable. The inclusion of $\lambda_{0}$ is crucial as it ensures trace preservation ($\operatorname{Tr}\rho=1$) and allows for the representation of any $3\times 3$ Hermitian operator, such as the density matrix $\rho$, Hamiltonian $H$, or interaction potential $V_{k}$, as a linear combination of these basis matrices as outlined in Eq. (3.6), Eq. (3.7) and Eq. (3.8) respectively. Notably, since $\lambda_{0}$ commutes with all $\lambda_{i}$, it does not interfere with the existing commutation relations, maintaining the necessary structure for neutrino dynamics analysis. Substituting above equations into the GKSL master Eq. (3.3) we obtain:
@@
$D_{\mu\nu}=\begin{pmatrix}-\Gamma_{00}&\beta_{01}&\beta_{02}&\beta_{03}&\beta_{04}&\beta_{05}&\beta_{06}&\beta_{07}&\beta_{08}\\ \beta_{01}&-\Gamma_{11}&\beta_{12}&\beta_{13}&\beta_{14}&\beta_{15}&\beta_{16}&\beta_{17}&\beta_{18}\\ \beta_{02}&\beta_{12}&-\Gamma_{22}&\beta_{23}&\beta_{24}&\beta_{25}&\beta_{26}&\beta_{27}&\beta_{28}\\ \beta_{03}&\beta_{13}&\beta_{23}&-\Gamma_{33}&\beta_{34}&\beta_{35}&\beta_{36}&\beta_{37}&\beta_{38}\\ \beta_{04}&\beta_{14}&\beta_{24}&\beta_{34}&-\Gamma_{44}&\beta_{45}&\beta_{46}&\beta_{47}&\beta_{48}\\ \beta_{05}&\beta_{15}&\beta_{25}&\beta_{35}&\beta_{45}&-\Gamma_{55}&\beta_{56}&\beta_{57}&\beta_{58}\\ \beta_{06}&\beta_{16}&\beta_{26}&\beta_{36}&\beta_{46}&\beta_{56}&-\Gamma_{66}&\beta_{67}&\beta_{68}\\ \beta_{07}&\beta_{17}&\beta_{27}&\beta_{37}&\beta_{47}&\beta_{57}&\beta_{67}&-\Gamma_{77}&\beta_{78}\\ \beta_{08}&\beta_{18}&\beta_{28}&\beta_{38}&\beta_{48}&\beta_{58}&\beta_{68}&\beta_{78}&-\Gamma_{88}\end{pmatrix}$. (3.10)
@@
In theory, the possibility of energy non-conservation within the neutrino subsystem exists, although the energy is inherently conserved for the entire system. However, we strictly adhere to the energy conservation condition within the neutrino subsystem for our analysis. Apart from this, conditions like entropy increase and the complete positivity of the neutrino density matrix further simplify the dissipator matrix. For simplicity, in this work, we consider a diagonal dissipator of the form as,
@@ 4.2 MDR-Induced Master Equation
It is to be noted here that the above equation denotes the unitary evolution of the entire system, i.e., the neutrino and the environment. However, in this study, we are interested in finding the decoherence of the neutrino sub-system that originates from the effective non-unitarity induced by the MDR. Therefore, to solve this equation for the density matrix neutrino subsystem, we move to the interaction picture, where the eigenstate which was previously in the Schrödinger picture, now takes the form,
@@
Note that from here onward the prime represents the respective quantities in the interaction picture. However, since $H_{\mathrm{I}},H_{0}]=0$, one can understand that $H_{\mathrm{I}}=H^{\prime}_{\mathrm{I}}$. A precise solution to this equation can be obtained by integrating both sides over the interval $[0,t]$, yielding,
$\int_{0}^{t}\dot{\rho}^{\prime}(t^{\prime})\,dt^{\prime}=-i\int_{0}^{t}\left[H_{\mathrm{I}},\rho^{\prime}(t^{\prime})\right]dt^{\prime},$ (4.15)
@@
$\rho^{\prime}(t)=\rho^{\prime}(0)-i\int_{0}^{t}\left[H_{\mathrm{I}},\rho^{\prime}(t^{\prime})\right]dt^{\prime}.$ (4.16)
@@
$\dot{\rho}^{\prime}(t)=-i\left[H_{\mathrm{I}},\rho^{\prime}(0)\right]-\int_{0}^{t}\left[H_{\mathrm{I}},\left[H_{\mathrm{I}},\rho^{\prime}(t^{\prime})\right]\right]dt^{\prime}.$ (4.17)
@@
$\dot{\rho}^{\prime}(t)=-i\left[H_{\mathrm{I}},\rho^{\prime}(0)\right]-\int_{0}^{t}\left[H_{\mathrm{I}},\left[H_{\mathrm{I}},\rho^{\prime}(t)\right]\right]dt^{\prime}.$ (4.18)
At this stage, we note that Quantum Gravity (QG) theories must include space-time fluctuations at the Planck scale. This can be done by introducing a fluctuating deformation parameter to address the fluctuating minimum length scale, requiring a stochastic treatment to capture the complexities involved. At this point, while averaging the fluctuations is a viable option, we have chosen not to pursue that approach in this study. Since the first term on the r.h.s. of Eq. (4.18) is commutator consists of time independent entities, it generates a global phase term in the neutrino oscillation probabilities and can therefore be neglected. Substituting $H_{\mathrm{I}}$ in terms of $H_{\mathrm{0}}$ from Eq. (4.8) in Eq. (4.18) in the second term, we get,
$\dot{\rho}^{\prime}(t)=-\kappa\left[H_{0},\left[H_{0},\rho^{\prime}(t)\right]\right],$ (4.19)
@@
This is the new master equation, which incorporates the decoherence effects arising because of the modified dispersion relation. It should be noted here the $\rho$ in the above above expression is the density matrix of the neutrino sub-system with the Born-Markov approximation in the Schrödinger representation and hence should not be confused with the density matrix of the total closed system expressed in Eq. (4.11). From Eq. (4.21), we can get nine coupled differential equations for neutrino density matrix as given by,
$\dot{\rho}_{0}(t)$ $=0\,$, (4.22a)
$\dot{\rho}_{1}(t)$ $=\frac{\Delta m^{2}_{21}}{2E}\rho_{2}(t)-\frac{(\Delta m^{2}_{21})^{2}t}{E^{2-2a}E_{P}^{2a}}\rho_{1}(t)\,$, (4.22b)
$\dot{\rho}_{2}(t)$ $=-\frac{\Delta m^{2}_{21}}{2E}\rho_{1}(t)-\frac{(\Delta m^{2}_{21})^{2}t}{E^{2-2a}E_{P}^{2a}}\rho_{2}(t)\,$, (4.22c)
$\dot{\rho}_{3}(t)$ $=0\,$, (4.22d)
$\dot{\rho}_{4}(t)$ $=\frac{\Delta m^{2}_{31}}{2E}\rho_{5}(t)-\frac{(\Delta m^{2}_{31})^{2}t}{E^{2-2a}E_{P}^{2a}}\rho_{4}(t)\,$, (4.22e)
$\dot{\rho}_{5}(t)$ $=-\frac{\Delta m^{2}_{31}}{2E}\rho_{4}(t)-\frac{(\Delta m^{2}_{31})^{2}t}{E^{2-2a}E_{P}^{2a}}\rho_{5}(t)\,$, (4.22f)
$\dot{\rho}_{6}(t)$ $=\frac{\Delta m^{2}_{32}}{2E}\rho_{7}(t)-\frac{(\Delta m^{2}_{32})^{2}t}{E^{2-2a}E_{P}^{2a}}\rho_{6}(t)\,$, (4.22g)
$\dot{\rho}_{7}(t)$ $=-\frac{\Delta m^{2}_{32}}{2E}\rho_{6}(t)-\frac{(\Delta m^{2}_{32})^{2}t}{E^{2-2a}E_{P}^{2a}}\rho_{7}(t)\,$, (4.22h)
$\dot{\rho}_{8}(t)$ $=0\,.$ (4.22i)
The r.h.s. of Eqs. (4.22a), (4.22d), and (4.22i) vanishes due to the specific form of the disspator chosen in Eq. (3.11). Furthermore, it can be seen from Eq. (4.21) that it takes different forms depending on the value of $a$, which is an input from the specific theory of quantum gravity that causes the modification to the dispersion relation. In this study, we remain agnostic about any such theories of quantum gravity. Later we are going to consider a few specific values of $a$ and compare their effects. At this point, we note that, the first terms on r.h.s. of Eq. (4.21) and Eqs. (4.22) account for trivial unitary dynamics of neutrino oscillations while the second terms govern neutrino-environment interactions due to MDR [^3] In the second term of the r.h.s of Eq. (4.21), the variable $t$ will denote the propagation time of the neutrinos. Since in this work we are mostly dealing with ultra-relativistic neutrinos, the propagation time $t$ can be replaced with the propagation length $L$ as shown in Eq. (6.1).. Moreover, the second terms on the r.h.s. of Eq. (4.21) and Eqs. (4.22) give insights about the form of decoherence strength parameter. Its explicit time-dependence comes through the dependence on $\rho$ itself which is further discussed in Sec. 6.2.